# Sample complexity of optimal transport barycenters with discrete support

### v2

@@ C.2 Sliced Wasserstein barycenter
then for all $\epsilon>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$:
$\mathcal{N}(\tau,\mathcal{F}_{p}^{s},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{12pR^{p}}{\tau}\right)^{d}\left(1+\frac{6pR^{p}}{\tau}\right)^{dN}\left(\frac{12pR^{p}}{\tau}\right)^{N}.$
Moreover $\mathcal{F}_{p,\mathbb{Q}}^{\text{s}}$, the subset of functions of $\mathcal{F}_{p}^{\text{s}}$ whose parameters $(Y,w,\theta)$ are rationals, is dense in $\mathcal{F}_{p}^{\text{s}}$.
@@
$\begin{split}\left|\ |\langle x-y_{i},\theta\rangle|^{p}-|\langle x-z_{i},\varphi\rangle|^{p}\right|&\leq pR^{p-1}|\langle x-y_{i},\theta\rangle-\langle x-z_{i},\varphi\rangle|\\ &=pR^{p-1}|\langle x,\theta-\varphi\rangle+\langle z_{i},\varphi-\theta\rangle+\langle z_{i},\theta\rangle-\langle y_{i},\theta\rangle|\\ &\leq pR^{p-1}(|\langle x,\theta-\varphi\rangle|+|\langle z_{i},\varphi-\theta\rangle|+|\langle z_{i}-y_{i},\theta\rangle|)\\ &\leq pR^{p-1}(2R\|\theta-\varphi\|+\|z_{i}-y_{i}\|).\end{split}$
@@
Theorem C.4.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Then for $\mathbb{D}=SW_{p}^{p}$ or $\text{max-}SW_{p}^{p}$ and for some integer $N\geq 1$, we have
$\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right]\leq C_{p,R}\sqrt{\frac{N(d+1)+d}{n}}.$
where $C_{p,R}=8\sqrt{2}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{48pR^{p}}{\tau}\right)}d\tau$.
Proof.
@@
For the sake of brevity, we only prove the result for $\mathbb{D}=SW_{p}^{p}$ as the $\text{max-}SW_{p}^{p}$ case is very similar. Let $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{B}_{R}^{M}\setminus D_{M})\times\mathrm{int}(\Delta_{N})$ be as in Proposition A.1. For all $\ell\in[\![1,L]\!]$, let $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ be i.i.d random variables of law $\mu^{\ell}$. With a slight abuse of notation and following the construction in Section A.3, we define the function $M:\theta\mapsto\operatorname{Card}(\{i\in[\![1,M]\!],\>\langle y_{i}-y_{j},\theta\rangle\neq 0,\forall j\neq i\})$. Then using Proposition B.1 and defining $K:=2pR^{p}$, we have:
@@
$\begin{split}&\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\tilde{\pi}_{i}\delta_{\tilde{y_{i}}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\tilde{\pi}_{i}\delta_{\tilde{y_{i}}}\right)\right|\\ &=\left|\frac{1}{L}\sum_{\ell=1}^{L}\left[\int_{\mathbb{S}^{d-1}}\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}}\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle X_{k}^{\ell}-\tilde{y}_{i},\theta\rangle|^{p}-w_{i}\}d\sigma(\theta)+\sum_{i=1}^{M(\theta)}\tilde{\pi}_{i}w_{i}\right.\right.\\ &\left.\left.-\int_{\mathbb{S}^{d-1}}\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle x-\tilde{y}_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)d\sigma(\theta)+\sum_{i=1}^{M(\theta)}\tilde{\pi}_{i}w_{i}\right]\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{\theta\in\mathbb{S}^{d-1}}{\max}{\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle X_{k}^{\ell}-y_{i},\theta\rangle|^{p}-w_{i}\}\right.\\ &\left.-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)\right|.\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{\theta\in\mathbb{S}^{d-1}}{\max}\underset{w_{i}=w_{j}\>\text{if}\>\langle y_{i}-y_{j},\theta\rangle=0}{\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{N}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,N}{\min}\{|\langle X_{k}^{\ell}-y_{i},\theta\rangle|^{p}-w_{i}\}\right.\\ &\left.-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,N}{\min}\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)\right|,\end{split}$
@@
where the first inequality uses Lemma D.2 for the maximum with respect to $w$, and then upper bound the integral with respect to $\theta$ by its maximum value. The last equality is due to the fact that for all $\theta\in\mathbb{S}^{d-1}$, $Y\in(\mathbb{R}^{d})^{N}$ and $x\in\mathbb{R}^{d}$, we have that
$\min_{k=1,\ldots,M(\theta)}\mid\langle x-y_{k},\theta\rangle\mid^{p}-w_{k}=\min_{k=1,\ldots,N}\mid\langle x-y_{k},\theta\rangle\mid^{p}-w_{k}$
whenever $w_{i}=w_{j}$ if $\langle y_{i}-y_{j},\theta\rangle=0$.
@@
$\begin{split}&\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\left|\operatorname{F}_{p}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{Y_{i}}\right)-\operatorname{F}_{p}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{Y_{i}}\right)\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\ \underset{(\theta,Y,w)\in\mathbb{S}^{d-1}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}}{\sup}\left|\mathbb{P}_{n}^{\ell}\left(f_{\theta,Y,w}\right)-\mathbb{E}_{X\sim\mu^{\ell}}\left[f_{\theta,Y,w}(X)\right]\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}^{\text{s}}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|,\end{split}$
@@
With the same argument as in the proof of Theorem C.2 this quantity can be upper bounded in expectation using (gine2021mathematical, Theorem 3.5.1):
@@ C.3 Entropy regularized barycenter
then for all $\tau>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$, we have
$\mathcal{N}(\tau,\mathcal{F}_{p}^{\epsilon},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{4pR^{p}}{\tau}\right)^{dN}\left(\frac{16p(2R)^{p}}{\tau}\right)^{N}.\\$
Moreover $\mathcal{F}_{p,\mathbb{Q}}^{\epsilon}$, the subset of functions of $\mathcal{F}_{p}^{\epsilon}$ whose parameters $(Y,\pi,w)$ are rationals, is dense in $\mathcal{F}_{p}^{\epsilon}$.
@@
$\begin{split}|g_{Y,\pi,w}^{\epsilon}(x)-g_{Z,\gamma,v}^{\epsilon}(x)|&=\epsilon\left|\log\left(\frac{\sum_{i=1}^{N}\pi_{i}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)}{\sum_{i=1}^{N}\gamma_{i}\exp\left(\frac{v_{i}-\|x-z_{i}\|^{p}}{\epsilon}\right)}\right)\right|\\ &\leq\epsilon\left|\log\left(\frac{\max_{i=1,\ldots,N}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)\sum_{i=1}^{N}\pi_{i}}{\min_{i=1,\ldots,N}\exp\left(\frac{v_{i}-\|x-z_{i}\|^{p}}{\epsilon}\right)\sum_{i=1}^{N}\gamma_{i}}\right)\right|\\ &=\left|\max_{i=1,\ldots,N}w_{i}-\|x-y_{i}\|^{p}-\min_{i=1,\ldots,N}v_{i}-\|x-z_{i}\|^{p}\right|\\ &\leq|w_{\overline{i}}-v_{\underline{i}}|+|\|x-z_{\underline{i}}\|^{p}-\|x-y_{\overline{i}}\|^{p}|\end{split}$
where $\overline{i}$ (resp. $\underline{i}$) is an index for which the maximum (resp. minimum) of $i\mapsto w_{i}-\|x-y_{i}\|^{p}$ (resp. $i\mapsto v_{i}-\|x-z_{i}\|^{p}$) is attained. By Lemma D.1, we thus have
@@
Notice that the right-hand side of the above inequality depends on neither $\pi$ nor $\epsilon$. Let $T=pR^{p-1}$, with the same proof argument of Theorem C.2, and by boundedness of the entropic dual variables (Proposition B.1) we thus get
@@
Theorem C.6.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let $\mathbb{D}=W_{\epsilon,p}^{p}$ with $\epsilon>0$, then for some integer $N\geq 1$:
$\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right]\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}}$
where $C_{p,R}=8\sqrt{2}\int_{0}^{(4p+1)(2R)^{p}}\sqrt{\log\left(2+\frac{64p(2R)^{p}}{\tau}\right)}d\tau$.
Proof.
We fix $\epsilon>0$. Let $(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{B}_{R}^{M}\setminus D_{M})\times\mathrm{int}(\Delta_{M})$ be as in Proposition A.1. For any triplet $(Y,\pi,w)\in\mathbb{B}_{R}^{N}\times\Delta_{N}\times\mathbb{R}^{N}$ and $x\in\mathbb{B}_{R}$, we define the function
$g_{Y,\pi,w}^{\epsilon}(x):=-\epsilon\log\left(\sum_{j=1}^{N}\pi_{j}\exp\left(\frac{w_{j}-\|x-y_{j}\|^{p}}{\epsilon}\right)\right).$
For all $\ell\in[\![1,L]\!]$, let $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ be i.i.d random variables of law $\mu^{\ell}$. Using Proposition B.1 and the same argument as in the proof of Theorem C.2, we get for $K:=4p(2R)^{p}$
@@
Then, by Lemma C.5, the bound obtained in (C.6) and with the same argument as in Theorem C.2, we obtain by (gine2021mathematical, Theorem 3.5.1) the following bound
@@
Corollary C.7.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be an integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $\mathbb{D}=\overline{W}_{\epsilon,p}^{p}$ (for some $\epsilon>0$). We denote by $(Y^{n},\pi^{n})$ an argmin in $A$ of
$(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right).$
Then we have
$\begin{split}&\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\\ &\leq 2C_{p,R}\sqrt{\frac{N(d+1)}{n}}+2K_{p,R}\left(\frac{e^{\frac{{\mathcal{K}}}{\epsilon}}}{\sqrt{n}}\left(1+\frac{1}{\epsilon^{\lfloor d/2\rfloor}}\right)\right),\end{split}$
with $\kappa=2C_{p}R+(2R)^{p}$, where $C_{p}=\sqrt{2}p(2R)^{p-1}$ is the Lipshitz constant of the Euclidean cost at the power $p$. Additionally, $K_{p,R}$ is an unknown constant depending only on $R$ and $p$ and $C_{p,R}$ is the constant associated to $W_{\epsilon,p}^{p}$ in Theorem 4.1.
Proof.
Let $(Y^{*},\pi^{*})$ and $(Y^{n},\pi^{n})$ be respective minimizers of $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ and $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$. We then denote $\nu^{*}=\sum_{i=1}^{N}\pi_{i}^{*}\delta_{y_{i}^{*}}$ and $\nu^{n}=\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}$. Note that the minimizers exist by continuity of the functional on the compact set $A$. Notice now that
@@
In the right-hand side of the inequality, we bound the first term by Theorem 4.1 while the second term can be bounded thanks to the proof of (genevay2019sample, Theorem 3). In fact, their proof can be adapted almost directly to show that
$\mathbb{E}|W_{\epsilon,p}^{p}(\mu_{n},\mu_{n})-W_{\epsilon,p}^{p}(\mu,\mu)|=O\left(\frac{e^{\frac{{\mathcal{K}}}{\epsilon}}}{\sqrt{n}}\left(1+\frac{1}{\epsilon^{\lfloor d/2\rfloor}}\right)\right).$
This result is given for arbitrary measures $\alpha,\beta\in\mathcal{M}_{1}(\mathbb{R}^{d})$, their empirical counterparts $\hat{\alpha}_{n},\>\hat{\beta}_{n}$ and the quantity $W_{\epsilon,p}^{p}(\hat{\alpha}_{n},\hat{\beta}_{n})$, but at no point in the proof the independence of the samples upon which $\hat{\alpha}_{n}$ and $\hat{\beta}_{n}$ are constructed is mandatory. As a result, one can obtain the same bound with $\beta=\alpha=\mu$.

### v3

@@ C.2 Sliced Wasserstein barycenter
then for all $\tau>0$, $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$, $\mathcal{N}(\tau,\mathcal{F}_{p}^{s},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{12pR^{p}}{\tau}\right)^{d}\left(1+\frac{6pR^{p}}{\tau}\right)^{dN}\left(\frac{12pR^{p}}{\tau}\right)^{N}.$ Also, the subset $\mathcal{F}_{p,\mathbb{Q}}^{\text{s}}$ of functions of $\mathcal{F}_{p}^{\text{s}}$ with rationals parameters $(Y,w,\theta)$, is dense in $\mathcal{F}_{p}^{\text{s}}$.
@@
$\begin{split}\left|\ |\langle x-y_{i},\theta\rangle|^{p}-|\langle x-z_{i},\varphi\rangle|^{p}\right|&\leq pR^{p-1}|\langle x-y_{i},\theta\rangle-\langle x-z_{i},\varphi\rangle|\\ &=pR^{p-1}|\langle x,\theta-\varphi\rangle+\langle z_{i},\varphi-\theta\rangle+\langle z_{i},\theta\rangle-\langle y_{i},\theta\rangle|\\ &\leq pR^{p-1}(2R\|\theta-\varphi\|+\|z_{i}-y_{i}\|).\end{split}$
@@
Proof of Theorem 3.1 for $\mathbb{D}=SW_{p}^{p}$ and $\text{max-}SW_{p}^{p}$.
@@
For brevity, we only prove the result for $\mathbb{D}=SW_{p}^{p}$ as the $\text{max-}SW_{p}^{p}$ case is similar. Let $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{B}_{R}^{M}\setminus D_{M})\times\mathrm{int}(\Delta_{N})$ be as in Proposition A.1. For all $\ell\in[\![1,L]\!]$, let $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ be i.i.d random variables of law $\mu^{\ell}$. Following the construction in Section A.3, we define the function $M:\theta\mapsto\operatorname{Card}(\{i\in[\![1,M]\!],\>\langle y_{i}-y_{j},\theta\rangle\neq 0,\forall j\neq i\})$. By Proposition B.1 and for $K:=2pR^{p}$, we get
@@
$\begin{split}&\text{Gen}_{n}^{\text{s}}(Y,\pi):=\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\tilde{\pi}_{i}\delta_{\tilde{y_{i}}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\tilde{\pi}_{i}\delta_{\tilde{y_{i}}}\right)\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{\theta\in\mathbb{S}^{d-1}}{\max}{\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle X_{k}^{\ell}-y_{i},\theta\rangle|^{p}-w_{i}\}\right.\\ &\left.-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)\right|.\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{\theta\in\mathbb{S}^{d-1}}{\max}\underset{w_{i}=w_{j}\>\text{if}\>\langle y_{i}-y_{j},\theta\rangle=0}{\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{N}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,N}{\min}\{|\langle X_{k}^{\ell}-y_{i},\theta\rangle|^{p}-w_{i}\}\right.\\ &\left.-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,N}{\min}\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)\right|,\end{split}$
@@
where the first inequality uses Lemma D.2 for the maximum with respect to $w$, and then upper bound the integral with respect to $\theta$ by its maximum value. The last equality is due to the fact that for all $\theta\in\mathbb{S}^{d-1}$, $Y\in(\mathbb{R}^{d})^{N}$ and $x\in\mathbb{R}^{d}$, we have that $\min_{k=1,\ldots,M(\theta)}\mid\langle x-y_{k},\theta\rangle\mid^{p}-w_{k}=\min_{k=1,\ldots,N}\mid\langle x-y_{k},\theta\rangle\mid^{p}-w_{k}$ whenever $w_{i}=w_{j}$ if $\langle y_{i}-y_{j},\theta\rangle=0$.
@@
$\begin{split}\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\text{Gen}_{n}^{\text{s}}(Y,\pi)&\leq\frac{1}{L}\sum_{\ell=1}^{L}\ \underset{(\theta,Y,w)\in\mathbb{S}^{d-1}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}}{\sup}\left|\mathbb{P}_{n}^{\ell}\left(f_{\theta,Y,w}\right)-\mathbb{E}_{X\sim\mu^{\ell}}\left[f_{\theta,Y,w}(X)\right]\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}^{\text{s}}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|,\end{split}$
@@
With the same argument as in the proof of case $\mathbb{D}=W_{p}^{p}$, this quantity can be upper bounded in expectation using (gine2021mathematical, Theorem 3.5.1):
@@ C.3 Entropy regularized barycenter
then for all $\tau>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$, we have $\mathcal{N}(\tau,\mathcal{F}_{p}^{\epsilon},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{4pR^{p}}{\tau}\right)^{dN}\left(\frac{16p(2R)^{p}}{\tau}\right)^{N}.$ Also, the subset $\mathcal{F}_{p,\mathbb{Q}}^{\epsilon}$ of functions of $\mathcal{F}_{p}^{\epsilon}$ with rational parameters $(Y,\pi,w)$, is dense in $\mathcal{F}_{p}^{\epsilon}$.
@@
$\begin{split}|g_{Y,\pi,w}^{\epsilon}(x)-g_{Z,\gamma,v}^{\epsilon}(x)|&=\epsilon\left|\log\left(\frac{\sum_{i=1}^{N}\pi_{i}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)}{\sum_{i=1}^{N}\gamma_{i}\exp\left(\frac{v_{i}-\|x-z_{i}\|^{p}}{\epsilon}\right)}\right)\right|\\ &\leq\epsilon\left|\log\left(\frac{\max_{i=1,\ldots,N}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)\sum_{i=1}^{N}\pi_{i}}{\min_{i=1,\ldots,N}\exp\left(\frac{v_{i}-\|x-z_{i}\|^{p}}{\epsilon}\right)\sum_{i=1}^{N}\gamma_{i}}\right)\right|\\ &=\left|\max_{i=1,\ldots,N}w_{i}-\|x-y_{i}\|^{p}-\min_{i=1,\ldots,N}v_{i}-\|x-z_{i}\|^{p}\right|.\end{split}$
Denoting $\overline{i}$ (resp. $\underline{i}$) is an index for which the maximum (resp. minimum) of $i\mapsto w_{i}-\|x-y_{i}\|^{p}$ (resp. $i\mapsto v_{i}-\|x-z_{i}\|^{p}$) is attained, we get by Lemma D.1,
@@
Notice that the right-hand side of the above inequality depends on neither $\pi$ nor $\epsilon$. Let $T=pR^{p-1}$, with the same proof argument as of case $\mathbb{D}=W_{p}^{p}$ and by boundedness of the entropic dual variables (Proposition B.1) we thus get
@@
Proof of Theorem 3.1 for $\mathbb{D}=W_{\epsilon,p}^{p}$.
We fix $\epsilon>0$. Let $(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{B}_{R}^{M}\setminus D_{M})\times\mathrm{int}(\Delta_{M})$ be as in Proposition A.1. For any triplet $(Y,\pi,w)\in\mathbb{B}_{R}^{N}\times\Delta_{N}\times\mathbb{R}^{N}$ and $x\in\mathbb{B}_{R}$, we let $g_{Y,\pi,w}^{\epsilon}(x):=-\epsilon\log\left(\sum_{j=1}^{N}\pi_{j}\exp\left(\frac{w_{j}-\|x-y_{j}\|^{p}}{\epsilon}\right)\right).$ For all $\ell\in[\![1,L]\!]$, let $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ be i.i.d random variables of law $\mu^{\ell}$. Using Proposition B.1 and the same argument as in the case $\mathbb{D}=W_{p}^{p}$, we get for $K:=4p(2R)^{p}$:
@@
Then, by Lemma C.3, the bound obtained in (C.6) and, once again, with the same argument as in the case $\mathbb{D}=W_{p}^{p}$, we obtain by (gine2021mathematical, Theorem 3.5.1) the following bound
@@
Proof of Corollary 3.6.
Let $(Y^{*},\pi^{*})$ and $(Y^{n},\pi^{n})$ be respective minimizers of $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ and $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$, which exist by continuity on the compact set $A$. We denote $\nu^{*}=\sum_{i=1}^{N}\pi_{i}^{*}\delta_{y_{i}^{*}}$ and $\nu^{n}=\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}$. Then
@@
In the right-hand side of the inequality, we bound the first term by Theorem 3.1 while the second term can be bounded thanks to the proof of (genevay2019sample, Theorem 3). In fact, their proof can be adapted almost directly to show that $\mathbb{E}|W_{\epsilon,p}^{p}(\mu_{n},\mu_{n})-W_{\epsilon,p}^{p}(\mu,\mu)|=O\left(\frac{e^{\frac{{\mathcal{K}}}{\epsilon}}}{\sqrt{n}}\left(1+\frac{1}{\epsilon^{\lfloor d/2\rfloor}}\right)\right).$ This result is given for arbitrary measures $\alpha,\beta\in\mathcal{M}_{1}(\mathbb{R}^{d})$, their empirical counterparts $\hat{\alpha}_{n},\>\hat{\beta}_{n}$ and the quantity $W_{\epsilon,p}^{p}(\hat{\alpha}_{n},\hat{\beta}_{n})$, but at no point in the proof the independence of the samples upon which $\hat{\alpha}_{n}$ and $\hat{\beta}_{n}$ are constructed is mandatory. As a result, one can obtain the same bound with $\beta=\alpha=\mu$.